Could someone please show me how to create an equation of a nonlinear function and provide two imputs?
Answer by stanbon(75887)

Could someone please show me how to create an equation of a nonlinear function and provide two imputs?
---
y = x^2 is the simplest possible non-linear quadratic function.
---
If x = 0,y = 0^2 = 0
If x = 1, y = 1^2 = 1
--------
Use Google to find oodles of information about non-linear functions.
================
Cheers,
Stan H.